OpenStudy (anonymous):

a • b • 4 = 4 • a • b is an example of which algebraic property?

OpenStudy (anonymous):

associative property of multiplication :)

OpenStudy (anonymous):

thank yu oh so very much

OpenStudy (anonymous):

no problem :)

OpenStudy (phi):

associative property means you can group the multiplication in different ways: (a*b*c)= (a*b)*c = a*(b*c) where it is understood you do the multiplication in parenthesis first. commutative property allows you to switch the order: a*b*c = b*a*c for example

OpenStudy (anonymous):

oooooh thanks for explaining now i get it :D